OpenStudy (anonymous):

A rectangular display case houses a square pyramid. Both have the same square base measuring x inches on a side. The top of the case is 2 inches above the tip of the pyramid which has a height of h inches. Which expression represents the volume of the empty space inside the case?

OpenStudy (anonymous):

a. x^3 - x^2(h-2) b. x^2-1/3x^2h c. 1/3x2^h d. 2/3x^2h e. x^2(h+2) - 1/3x^2h

OpenStudy (anonymous):

DO YOU KNOW IT

OpenStudy (anonymous):

volume of the whole case = x^2(h+2) volume of the pyrmid = (1/3)x^2h so volume of the empty space is x^2(h+2) - (1/3)x^2h

OpenStudy (anonymous):

EDIT: wrote wrong answer I'm not sure, but I think its D. I think the volume if a square pyramid is 1/3 base * height, so the remaining volume would be the other 2/3's